OpenStudy (anonymous):

What is the slope of the line that passes through the points (3, –1) and (–2, –5)? A. -4/5 B. -5/4 C. 5/4 D. 4/5

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

what answer is it

OpenStudy (anonymous):

=(-5-(-1))/(-2-3)=(-5+1)/-5=-4/-5=4/5

OpenStudy (anonymous):

k